# Audio-Language Models for Audio-Centric Tasks: A survey

### v1

@@ III-B2 Generative Objectives
Generative methods have proven to be powerful and effective in audio representation learning. They lead the network in learning semantic features of audio through pretext tasks such as masked reconstruction [38]. In audio-language pre-training, similar approaches are introduced, guiding representation learning through audio or audio-related language generation tasks. These methods are often combined with contrastive learning to bolster the robustness of learned audio embeddings or improve computational efficiency. During transfer, these generative objectives can help the model adapt to corresponding generative tasks and are widely used in transfer with generative LLMs.
During pre-training, the most common method for audio mask reconstruction is based on the audio spectrogram. Let $M\left(\cdot\right)$ denote the masking operation, and let $f_{a}\left(\cdot\right)$ and $p_{ae}\left(\cdot\right)$ represent the spectrogram encoder and audio embedding projection layer, respectively. To achieve masked spectrogram prediction, an additional decoder $f_{a}^{-1}\left(\cdot\right)$ is added to the model. For an audio sample with the original spectrogram $a$, spectrogram reconstruction can be represented as $\hat{a}=f_{a}^{-1}(p_{ae}(f_{a}(M(a))))$. Using $\hat{a}_{n}$ and $a_{n}$ to denote the decoder prediction output of the $n-th$ masked spectrogram patch and the original true patch, respectively. For a spectrogram divided into $N$ patches, the audio reconstruction loss used for self-supervision can be defined as minimizing the $L2$ (mean squared error, MSE) loss:
@@
Since ALMs include both audio and language modalities as inputs, some works have similarly designed masked cross-modal reconstruction tasks, which typically involve methods such as cross-attention mechanisms to communicate between the encoders of the two modalities and perform reconstruction on the audio representation.
During audio generation transfer, training objectives essentially enhance the model’s performance by minimizing the distance between the predicted embedding $\hat{z}$ and its corresponding ground truth $z$. This distance metric can be chosen based on the situation, with common options including $L1$ and $L2$ distances. The training objective can also be set as a weighted sum of multiple distances. For an audio sample, generative audio modeling objective can be represented as:
@@
where $T$ denotes the total number of frames, $L$ denotes embedding dimension, and $\alpha$ and $\beta$ are weight hyperparameters. In addition to the method that uses embedding differences as a training objective, it is also possible to directly train jointly with the decoder network, designing the training objective directly on the predicted audio amplitude. For instance, aiming to learn a decoder net $h_{de}\left(\cdot\right)$ that maps known audio $x_{i}$ and query $t_{i}$ to a predicted audio $\hat{a}_{i}$. Denote $z_{i}^{t}$ as the embedding of the language query, the training objective could be to minimize the $L1$ (mean absolute error, MAE) loss between the amplitude spectrogram $|a_{i}|$ of the ground truth target audio source and the predicted $|\hat{a}_{i}|$:
$|\hat{a}_{i}|=h_{de}\left(z_{i}^{t}\right)$ (8)
$\mathcal{L}^{{}^{\prime}}_{am}=\sum_{i=1}^{B}\left\||{a}_{i}|-|\hat{a}_{i}|\right\|_{1}$ (9)
Generative language modeling objectives are used to guide ALM in generating audio-related text that is consistent with the ground truth. On one hand, they can be used to force the model to learn audio-language correlations to promote representation learning, and help improve the model’s performance on corresponding downstream tasks (e.g., automatic caption generation). On the other hand, as a standard loss for generative language modeling, it is also commonly used during ALM transfer with language model [39].
An additional text decoder (language pre-trained model) is required in language generation. When using an autoregressive language model to predict tokenized text associated with a given audio sample $x$, the language modeling objective is defined as minimizing the negative log-likelihood of the current ground-truth token (cross-entropy, CE loss), given the previous ground-truth tokens:
@@
Here, $y_{t}$ is the $t-th$ ground-truth token of the given caption $y$, $T$ is the total length of the caption, and $\theta$ represents the model’s learnable parameters. Non-autoregressive language models also adopt a similar negative log likelihood objective without temporal averaging.
@@ III-B3 Discriminative Objectives
where $C$ is the number of classes. $y_{i,c}$ is the true label of the $i-th$ sample in class $c$ (0 or 1). $\hat{p}_{i,c}$ is the predicted probability of the $i-th$ sample in class $c$.
Audio-Text Retrieval (ATR) aims to find matching items between audio clips and textual descriptions. Given a query in one modality (audio or text), the goal is to retrieve the corresponding item from a pool of candidates in the other modality. Here, we use a scoring function $S\left(\cdot\right)$ to represent the model’s prediction output by measuring the correlation between audio and text. Denote $Y$ as a set of $m$ possible caption texts, the correspondence caption of a given audio $x_{i}$ is
@@
Specially, audio-text matching is pretext task designed to forcing a more fine-grained alignment between audio and text embeddings than contrastive pre-training. It train the model to predict whether a given text correctly describes a provided audio, can be seen as a binary classification task requiring the model to determine whether an audio-language pair is a match or not. The matching objective can be defined as:
@@
III-C Evaluation Methods
Model evaluation aims to fairly measure the performance of models under the same experimental setup and tasks. The evaluation methods for ALMs mainly include zero-shot (ZS), linear probe, supervised fine-tuning, and instruction-following evaluation. Each of these methods has their own focus, collectively forming the basis for a comprehensive performance evaluation of ALMs.
III-C1 Zero-Shot Evaluation
It focuses on assessing the ability of contrastive ALMs in open-set retrieval. This zero-shot prediction is primarily conducted by measuring the similarity between audio and text embeddings. Notably, aside from direct text-to-audio or audio-to-text retrieval, considering that labels are also a special form of language. This allows for zero-shot evaluation on classification tasks such as sound event detection and emotion recognition.
III-C2 Linear Probe Evaluation
It is a common experimental setup for evaluating pre-trained models, and it is used to assess the audio representation of ALMs. It involves adding a linear header (usually an MLP) to the frozen pre-trained model and training the header on downstream tasks, allowing the model to be adapted for specific tasks and datasets. Although this simple transfer learning setup may not achieve optimal performance on specific tasks, it minimizes the variables introduced, hence its widespread adoption for conducting fair representational evaluations. In linear probe evaluation, the selected tasks are usually fundamental linear tasks like classification.
III-C3 Supervised Fine-tune Evaluation
It further examines the generalization ability of the pre-trained model to downstream tasks and its task-specific performance. For a given downstream task, the audio encoder is unfrozen and fine-tuned along with an attached head. The model’s performance is then validated on the test set and compared with state-of-the-art (SOTA) models for that task. This evaluation approach not only covers classification tasks but often combines various types of tasks including detection and generation, thoroughly testing the model’s adaptability and flexibility.
III-C4 Instruction-following Evaluation
It emphasizes the model’s ability to understand complex instructions and respond accurately, serving as an important indicator of the task generalization ability of LALMs. This evaluation method can be considered a special type of zero-shot evaluation or supervised fine-tuning evaluation, depending on whether instruction tuning is performed.
@@ IV Representation Pre-training
Audio encoders and text encoders are the most critical components of ALMs. They provide the initialization of model parameters for post pre-training on audio-text pairs or transfer on downstream task datasets. Recent studies have shown that the choice of encoders significantly impacts the generation of powerful representations through Audio-language pre-training and enhances the performance of downstream tasks [40].
@@ IV-A Audio Pre-training
From the perspective of model architecture, pre-trained models used to initialize encoder in ALM mainly include convolutional neural network (CNN)-based, Transformer-based, and codec-based models.
IV-A1 CNN-based Models
CNNs are widely used in audio pre-training due to their strong feature extraction capabilities, parameter sharing, and sparse connectivity. In audio applications, CNNs-based models usually use short-time Fourier transform (STFTs) to convert time-domain waveforms into log-mel spectrograms as input. Models like AlexNet and VGG perform well in audio classification tasks[41, 42], and their performance is related to the design of network depth and width. PANNs [43] based on CNN14 achieve good results in Audioset labeling, highlighting the effectiveness of CNNs in learning audio representations.
IV-A2 Transformer-based Models
These models utilize self-attention mechanisms to capture long-range dependencies in audio signals and enhance the understanding of complex patterns. Wav2vec 2.0 [44] and HuBERT [45] are designed for speech tasks, combining CNN with self-attention to manage local features and long-term temporal dynamics. Wav2vec 2.0 uses a unique self-supervised method, which converts raw audio into potential representations through CNN, and then generates contextual output with masks and Transformer. Training involves contrastive tasks to predict masked segments to enhance generalization ability. HuBERT improves this by introducing noise labels from offline clustering to form a codebook for pre-text prediction, encouraging the model to capture acoustic details and long-range temporal relationships. Considering supervised models face overfitting risks during fine-tuning [46], Whisper [47] addresses this issue with a multi-task training framework, which uses an encoder-decoder Transformer with additional convolutional layers for processing audio spectrograms, promoting next token prediction across tasks and offering wide adaptability. AST [48] is the first model without convolutions and based purely on attention for audio spectrograms. While flexible, it requires more data for training and faces challenges with large GPU memory and long training times [49]. HTSAT [50] deals with these issues through a hierarchical model structure, processing audio signals through multiple layers, with each layer’s Transformer capturing temporal and structural information, enabling more effective handling of long audio signals. AudioMAE [38] explores the audio extension of the Masked Autoencoder (MAE) [51], adopting a Transformer-based encoder-decoder design, masking log-mel spectrograms at a high rate, with the encoder providing non-masked tokens and the decoder reconstructing masked parts. After fine-tuning, it achieves advanced performance in classification tasks and is one of the mainstream self-supervised pre-training models in the audio domain.
IV-A3 Codec-based Models
Based on the encoder-decoder structure, they can convert continuous audio into discrete tokens for developing ALMs. Despite some loss of information, they excel at extracting acoustic features and achieving high-quality audio reconstruction [52]. SoundStream [53] is a pioneer, using streamable SEANets [54] for encoding and decoding, and Residual Vector Quantization (RVQ) bottleneck for parallel token flow, optimized through reconstruction and adversarial loss. Encodec [55] enhances this framework by adding LSTM layers and using a Transformer-based language model to improve sequence modeling performance.
@@ IV-B Language Pre-training
Language modeling is a fundamental area in artificial intelligence. Based on parameter scale, language pre-training models can be categorized into small ($\leq 1$ billion), medium ($1-10$ billion), large ($10-100$ billion), and very large ($>100$ billion). Models with over $1$ billion parameters are referred to as LLMs. By pre-training on vast amounts of text data, these models learn rich semantic representations and grammatical structures. With the development of fine-tuning techniques incorporating human feedback, numerous LLMs with tens of billions of parameters have emerged, demonstrating promising capabilities for multimodal tasks. Renowned language pre-training models used as text encoders in ALMs including: GPT (Generative Pre-trained Transformers) families [56, 57, 58, 59, 60, 61], LLaMA (Large Language Model Meta AI) families [62, 63, 64, 65], Qwen [66, 67, 68, 69], and OPT (Open Pre-trained Transformer) families [70, 71]. Note that language pre-training is not the main focus of this paper, for more details please refer to [11, 72].
@@ IV-C Audio-Language Pre-training
Model Pre-trained Models Objectives Main Contribution
@@
MS-CLAP [2] CNN14+BERT first to propose contrastive learning for ALM pre-training.
LAION-CLAP [22] HTSAT+RoBERTa first open-source contrastive learning for ALM pre-training.
MS-CLAP V2 [50] HTSAT-22+GPT-2 employ a multi-task audio encoder and an autoregressive text encoder.
BLAT [73] CNN14+BERT utilize audio captioning to synthetic data for pre-training.
ACBA [74] CNN14+RoBERTa propose to synthesize sequential audio with captions to enhance sequential relationship modeling capability between sound events.
COMPA [26] HTSAT+Flan-T5 propose the composition-aware contrastive learning to enhance compositional reasoning capability.
MGA-CLAP [75] HTSAT/AST+BERT develop a shared codebook and hard negative guided contrastive loss to unify the granularity and latent distribution of two modalities.
MusCALL [76] ResNet50+GPT-2 excels strong performance in zero-shot genre classification and auto-tagging tasks within the music domain.
CTAL [28] RoBERTa early exploration of joint representation learning methods for audio and language modalities.
FLAP [77] MAViL+RoBERTa learning self-supervised audio-language representations by employing masking for both contrastive learning and reconstruction.
M2D-CLAP [78] M2D+GTE combine Masked Modeling Duo (M2D) to train audio encoder while contrastive learning.
Cacophony [79] AudioMAE+RoBERTa integrate auxiliary captioning to enhance audio encoding for fine-grained pattern matching with text descriptions.
MINT [34] Data2vec+Flan-T5 boost audio-language models through multi-target pre-training and instruction tuning.
  • Con: Contrastive objectives, Gen: Generative objectives, Dis: Discriminative objectives
@@ IV-C1 Contrastive Pre-trained Models
Contrastive learning is a common pre-training method that learns feature representations by distinguishing similar sample pairs (positive pairs) from dissimilar ones (negative pairs). Inspired by the CLIP [80] model, which leverages large-scale web visual-text pairs to expand the training scale for visual tasks and achieves excellent performance across various tasks, researchers have begun to apply this contrastive training paradigm to the audio field. Microsoft’s MS-CLAP [2] (prefix added by the editor for distinction) is the first contrastive language-audio pre-training model. It uses the symmetric audio-text infoNCE loss function 5 for pre-training based on audio-text paired datasets and audios from other tasks such as audio classification.
Subsequent studies have focused on dataset scaling. LAION-CLAP [22] released a larger audio-caption dataset, LAION-Audio-630K, and trained the first fully open-source CLAP model based on this dataset together with other public datasets. MS-CLAP V2 [50] not only leverages a more extensive multi-task trained audio encoder but also further expands the audio-text paired dataset for contrastive language-audio learning. BLAT [73], from another prospective, proposes using an audio captioning model to generate audio-text pair data for contrastive pretraining.
Another line of research aims to address the inherent shortcomings of the vanilla CLAP. The experiments of ACBA [74] show that CLAP has limited understanding of natural language, especially regarding the order or concurrent arrangement of sound events. It suggests modifying the original pre-training dataset to provide more audio-language pairs about ordering. COMPA [26] addresses the issue that current benchmarks cannot measure the lack of combinatorial reasoning in models and proposes contrastive training with composition-aware hard negatives and a modular contrastive loss to improve combinatorial reasoning capabilities. MGA-CLAP [75] tackles the problem of different granularities between modalities by adopting a shared codebook, designing a locality-aware block and a hard-negative guided loss to achieve fine-grained alignment.
Additionally, MusCALL [76] conducts pre-training research in professional fields such as music and performs excellently in related tasks.
@@ IV-C2 Generative Pre-trained Models
Generative pre-training aims to learn deeper semantic representations by setting generative audio or text as pretext tasks. Cross-modal Transformer for Audio-and-Language (CTAL) [28] is an early exploration of audio-language pretraining through masked language modeling and cross-attention based masked cross-modal acoustic modeling. Fast Language-Audio Pre-training (FLAP) [77] conduct representation learning through the combination of masked reconstruction and contrastive learning. It generates multiple augmented views of the audio through masking for inter-modal contrast and learns to reconstruct the masked parts of the audio spectrogram. This masking reduces the amount of data that needs to be processed, thereby lowering the computational complexity and making it more efficient than contrastive learning with raw spectrograms. Additionally, by incorporating the masked reconstruction task, the model is encouraged to compress information into each sample embedding, making the audio embedding not only close to their textual counterparts but also producing more informative original inputs. M2D-CLAP [78] addresses the issue of MAE using all patches to encode training signals, which may lead to underutilization of inductive biases. By combining Masked Modeling Duo (M2D) [81] to train the audio encoder while contrastive learning further promotes input modeling, thus enhancing the effectiveness of the learned representations. Practice in the vision-language has shown that integrating auxiliary captioning objectives in contrastive learning can provide stronger supervision [82, 83]. Cacophony [79] improves CLAP by incorporating an auxiliary captioning objective, encouraging the audio encoder to capture fine-grained patterns closely matching text descriptions.

### v2

@@ III-B2 Generative Objectives
Generative methods effectively learn audio representations through pretext tasks like masked reconstruction of audio and associated text [huang2022audiomae]. These methods are often combined with contrastive learning, they enhance embedding robustness and computational efficiency. In transfer learning, these objectives adapt models to downstream generation tasks and are widely used in generative LLM scenarios.
During pre-training, the most common method for audio mask reconstruction operates on spectrograms. Let $M(\cdot)$ denote the masking operation, $f_{a}(\cdot)$ the spectrogram encoder, and $p_{ae}(\cdot)$ the audio embedding projection layer. Masked spectrogram prediction incorporates an additional decoder $f_{a}^{-1}(\cdot)$. For an original spectrogram $a$, the reconstruction is formulated as $\hat{a}=f_{a}^{-1}(p_{ae}(f_{a}(M(a))))$. Denoting the decoder’s prediction and the original true value for the $n$-th masked spectrogram patch as $\hat{a}_{n}$ and $a_{n}$ respectively, the self-supervised audio reconstruction loss for $N$ patches minimizes the $L2$ (mean squared error, MSE) loss:
@@
As ALMs incorporate both audio and language modalities, some works design masked cross-modal reconstruction tasks. These typically use cross-attention mechanisms to enable communication between the two encoders and reconstruct the audio representation.
When conducting transfer through audio generation tasks, the training minimizing the distance between the predicted embedding $\hat{z}$ and ground truth $z$. Common distance metrics include $L1$ and $L2$ norms, with the objective often formulated as their weighted sum:
@@
where $T$ is the total frames, $L$ the embedding dimension, and $\alpha$, $\beta$ are weight hyperparameters. Alternatively, the objective can be applied directly to audio amplitude. For a decoder network $h_{de}(\cdot)$ that maps input audio $x_{i}$ and text query $t_{i}$ to predicted audio $|\hat{a}_{i}|$, with $z_{i}^{t}$ as the language embedding, the $L1$ loss between amplitude spectrograms is:
$|\hat{a}_{i}|=h_{de}\left(z_{i}^{t}\right),\quad\mathcal{L}^{{}^{\prime}}_{am}=\sum_{i=1}^{B}\left\||{a}_{i}|-|\hat{a}_{i}|\right\|_{1}$ (7)
Generative language modeling objectives guide ALMs to produce audio-related text consistent with ground truth. These objectives help learn audio-language correlations for improved performance on downstream tasks (e.g., automatic captioning) and it serve as standard losses during ALM transfer with language models [JMLR2023palm].
Language generation requires additional text decoder (e.g., a pre-trained language model), where the autoregressive objective minimizes the negative log-likelihood of each ground-truth token given the preceding tokens and audio input $x$:
@@
Here $y_{t}$ is the $t$-th ground-truth token of caption $y$, $T$ is the caption length, and $\theta$ represents model parameters. Non-autoregressive models use a similar negative log-likelihood objective without temporal averaging.
@@ III-B3 Discriminative Objectives
where $C$ is the number of classes. $y_{i,c}$ is the true label of the ithi-th sample in class $c$ (0 or 1). $\hat{p}_{i,c}$ is the predicted probability of the ithi-th sample in class $c$.
Audio-Text Retrieval (ATR) aims to find matching items between audio clips and textual descriptions. Given a query in one modality (audio or text), it retrieves the corresponding item from candidate pools in the other modality. Here, we use a scoring function $S\left(\cdot\right)$ to represent the model’s prediction output by measuring the correlation between audio and text. Denote $Y$ as a set of $m$ possible caption texts, the correspondence caption of a given audio $x_{i}$ is
@@
Specifically, audio-text matching is a pretext task designed to force a more fine-grained alignment between audio and text embeddings than contrastive pre-training. This task trains the model to predict whether a given text correctly describes a provided audio segment, functioning as a binary classification task that determines whether an audio-language pair matches or not. The matching objective can be defined as:
@@ IV Representation Pre-training
Audio encoders and text encoders are the most critical components of ALMs. They provide the initialization of model parameters for post pre-training on audio-text pairs or transfer on downstream task datasets. Recent studies have shown that the choice of encoders significantly impacts the generation of powerful representations through Audio-language pre-training and enhances the performance of downstream tasks [mei2024wavcaps].
@@ IV-A Audio Pre-training
From the perspective of model architecture, pre-trained models used to initialize encoders in ALMs mainly include CNN-based, Transformer-based, and codec-based models. CNN-based models like PANNs [kong2020panns] extract features from spectrograms and are widely used in audio classification. Transformer-based models dominate current research, with representative works including Wav2vec 2.0 [baevski2020wav2vec], HuBERT [hsu2021hubert], Whisper [radford2023whisper], AST [gong2021ast], and AudioMAE [huang2022audiomae] for speech and general audio tasks. Codec-based models such as SoundStream [zeghidour2021soundstream] and Encodec [fossez2023high] discretize audio via encoder-decoder structures for language modeling. Due to space limitations, detailed introductions of these models are provided in Supplementary Material.
@@ IV-B Language Pre-training
Language pre-training provides foundational text encoders for ALMs, enabling them to process and generate natural language descriptions related to audio content. These encoders are typically initialized with pre-trained language models before being jointly optimized with audio encoders during audio-language pre-training. Popular families include GPT [radford2019GPT2, achiam2023gpt4], LLaMA [touvron2023llama2, chiang2023vicuna], Qwen [bai2023qwen], and OPT [zhang2022opt]. For comprehensive reviews of language pre-training, we refer readers to [zhao2023LLMsurvey, minaee2024LLMs].
@@ IV-C Audio-Language Pre-training
[…]
@@
MS-CLAP [elizalde2023clap] CNN14+BERT
LAION-CLAP [wu2023large] HTSAT+RoBERTa
MS-CLAP V2 [elizalde2024CLAP2] HTSAT-22+GPT-2
BLAT [xu2023blat] CNN14+BERT
ACBA [wu2023ACBA] CNN14+RoBERTa
COMPA [ghosh2024compa] HTSAT+Flan-T5
MGA-CLAP [li2024MGACLAP] HTSAT/AST+BERT
T-CLAP [yuan2024tclap] HTSAT+RoBERTa
MusCALL [manco2022contrastive] ResNet50+GPT-2
CTAL [li2021ctal] RoBERTa
FLAP [yeh2023flap] MAViL+RoBERTa
M2D-CLAP [niizumi2024m2dclap] M2D+GTE
Cacophony [zhu2024cacophony] AudioMAE+RoBERTa
MINT [zhao2024mint] Data2vec+Flan-T5
[…]
  • Objectives: Contrastive (Con), Generative (Gen), Discriminative (Dis).
@@ IV-C1 Contrastive Pre-trained Models
Contrastive learning is a common pre-training method that learns feature representations by distinguishing similar sample pairs (positive pairs) from dissimilar ones (negative pairs). Inspired by the CLIP [RadfordCLIP2021] model, which leverages large-scale web visual-text pairs to expand the training scale for visual tasks and achieves excellent performance across tasks, researchers have apply this contrastive training paradigm to the audio field. Microsoft’s MS-CLAP [elizalde2023clap] (prefix added by the editor for distinction) is the first contrastive language-audio pre-training model. It uses the symmetric audio-text infoNCE loss function 4 for pre-training based on audio-text paired datasets and audios from other tasks such as audio classification.
Subsequent studies have focused on dataset scaling. LAION-CLAP [wu2023large] released a larger audio-caption dataset, LAION-Audio-630K, and trained the first fully open-source CLAP model together with other public datasets. MS-CLAP V2 [elizalde2024CLAP2] not only leverages a more extensive multi-task trained audio encoder but also further expands the audio-text paired dataset for contrastive learning. BLAT [xu2023blat], from another prospective, proposes using an audio captioning model to generate audio-text pair data for contrastive pretraining.
Another line of research aims to address the inherent shortcomings of the vanilla CLAP. The experiments of ACBA [wu2023ACBA] show that CLAP has limited understanding of natural language, especially regarding the order or concurrent arrangement of sound events. It suggests modifying the original pre-training dataset to provide more audio-language pairs about ordering. COMPA [ghosh2024compa] addresses the issue that current benchmarks cannot measure the lack of combinatorial reasoning in models and proposes contrastive training with composition-aware hard negatives and a modular contrastive loss to improve combinatorial reasoning capabilities. MGA-CLAP [li2024MGACLAP] tackles the problem of different granularities between modalities by adopting a shared codebook, designing a locality-aware block and a hard-negative guided loss to achieve fine-grained alignment. T-CLAP [yuan2024tclap] identifies its weakness in capturing temporal information and introduces a temporal-contrastive loss, using LLM-generated ordered and misordered captions to learn the sequence of sound events.
Additionally, research has also been conducted in specialized fields like music, with works such as MusCALL [manco2022contrastive] achieving excellent performance on relevant tasks.
@@ IV-C2 Generative Pre-trained Models
Generative pre-training aims to learn deeper semantic representations by setting generative audio or text as pretext tasks. Cross-modal Transformer for Audio-and-Language (CTAL) [li2021ctal] is an early exploration of audio-language pretraining through masked language modeling and cross-attention based masked cross-modal acoustic modeling. Fast Language-Audio Pre-training (FLAP) [yeh2023flap] conduct representation learning through the combination of masked reconstruction and contrastive learning. It generates multiple augmented views of the audio through masking for inter-modal contrast and learns to reconstruct the masked parts of the audio spectrogram. This masking reduces the amount of data that needs to be processed, thereby lowering the computational complexity and making it more efficient than contrastive learning with raw spectrograms. Additionally, by incorporating the masked reconstruction task, the model is encouraged to compress information into each sample embedding, making the audio embedding not only close to their textual counterparts but also producing more informative original inputs. M2D-CLAP [niizumi2024m2dclap] addresses the issue of MAE using all patches to encode training signals, which may lead to underutilization of inductive biases. By combining Masked Modeling Duo (M2D) [niizumi2023m2d] to train the audio encoder while contrastive learning further promotes input modeling, thus enhancing the effectiveness of the learned representations. Practice in the vision-language has shown that integrating auxiliary captioning objectives in contrastive learning can provide stronger supervision [yu2022coca, li2022blip]. Cacophony [zhu2024cacophony] improves CLAP by incorporating an auxiliary captioning objective, encouraging the audio encoder to capture fine-grained patterns closely matching text descriptions.